OpenStudy (anonymous):

Simplify. Do not leave negative exponents in your answer. (x^9/(-27y^3))^(-2/3)

OpenStudy (anonymous):

\[(x^9/-27y^3)^-2/3\]

OpenStudy (pfenn1):

\[\left( \frac{x^9}{-27y^3} \right)^{-2/3}\]Recall that \[\large {x^{\frac ab}=\sqrt[b]{x^a}}\]and \[\large x^{-a}=\frac{1}{x^a}\]

OpenStudy (pfenn1):

\[\left( \frac{x^9}{-27y^3} \right)^{-2/3}=\sqrt[3]{\left( \frac{x^9}{-27y^3} \right)^{-2}}=\sqrt[3]{\left( \frac{-27y^3}{x^9} \right)^2}\]

OpenStudy (pfenn1):

I'm not completely sure about the negative sign.

OpenStudy (pfenn1):

Without the negative sign, I think this would simplify down to \[\pm \frac{9y^2}{x^6}\]

OpenStudy (pfenn1):

Maybe\[\pm \frac{9y^2}{x^6}i\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

= [ ( - 3³ y³ )/ x^9 ]^ (2/3) = ( -3y/x³ )² = 9y²/ x^6